Find another tangent to y=1-3x+12x^2-8x^3 which is parallel to the tangent at (1,2).
Answer by Alan3354(69443)

Find another tangent to y=1-3x+12x^2-8x^3 which is parallel to the tangent at (1,2).
==============
y=1-3x+12x^2-8x^3
Find the 1st derivative
y' = -3 + 24x - 24x^2
At x=1, y' = -3
----------------------
y' = -3 + 24x - 24x^2 = 3
x = 0, 1
==============
For x = 0, y = -3x +1 is parallel and tangent